OpenStudy (anonymous):

Mark is in training for a marathon race. He has been training by running 10 miles each day. He will be increasing his running by this weekly formula: y=10(1.1)x . Using the formula, what distance, in miles, will Mark be running in 10 weeks? Round to the nearest mile. a. 18 miles b. 20 miles c. 24 miles d. 26 miles

OpenStudy (anonymous):

@JamesJ : but then this would be 10 but thats not an option..

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

@zzr0ck3r @lgbasallote

OpenStudy (anonymous):

You wrote it wrong. \[y = 10(1.1)^{x}\] Try that now.

OpenStudy (anonymous):

i got 11 ?

OpenStudy (anonymous):

hello ?

OpenStudy (anonymous):

Nope. I got 25.9 y = 10(1.1)^10 y = 10(2.59374246)

OpenStudy (jamesj):

Right, here x = 10, not x = 0.

OpenStudy (jamesj):

Thus \( 1.1^x \neq 1 \)